# Solvable Tuple Patterns and Their Applications to Program Verification

### v1

@@ 2.2 Solvable Tuple Patterns
The name “solvable tuple patterns” comes from the following fact. Suppose $t=(p_{1},\ldots,p_{n})\in\mathcal{T}_{n}$. Then, the solution for a system of equations $p_{1}=s_{1},\ldots,p_{n}=s_{n}$ can be expressed by using (i) $s_{1},\ldots,s_{n},\epsilon$, (ii) constants $a\in\Sigma$, and (iii) the (partial) operation $\ss\backslash s^{\prime}$, which is defined as $s_{0}$ such that $ss_{0}=s^{\prime}$ if $s$ is a prefix of $s^{\prime}$. For example, $t=(x_{1}x_{2},x_{2}x_{1},x_{1})$ is solvable in the sense that the system of equations $x_{1}x_{2}=s_{1},x_{2}x_{1}=s_{2},x_{1}=s_{3}$ has a general solution $x_{1}=s_{3},x_{2}=s_{3}\backslash s_{1}$ provided $(s_{3}\backslash s_{1})\backslash s_{2}=s_{3}$ (or, equivalently $(s_{3}\backslash s_{1})\cdot s_{3}=s_{2}$). ∎
As stated in the following theorems, our algorithm indeed infers solvable patterns, and is complete for the class of solvable tuple patterns (modulo a minor condition).
@@
The completeness theorem above would not hold if the assumption $M\models_{s}t$ were weakened to $M\models t$. In fact, let $t=(x,x)$ and $M=(\epsilon\ \epsilon)$. Then $M\models t$ and $t\in\mathcal{T}_{2}$, but we can only obtain $(\epsilon,\epsilon),(\epsilon,x),(x,\epsilon),(x,y)$ by rewriting. In fact, the only possible reduction sequences are:
$((x_{1},x_{2}),[M/(x_{1},x_{2})])\longrightarrow((x_{1},\epsilon),[(\epsilon)/(x_{1})])\longrightarrow((\epsilon,\epsilon),[\,])$
$((x_{1},x_{2}),[M/(x_{1},x_{2})])\longrightarrow((\epsilon,x_{2}),[(\epsilon)/(x_{2})])\longrightarrow((\epsilon,\epsilon),[\,]).$
The rule R-Prefix is inapplicable because of the side condition $M[*][i]\neq\widetilde{\epsilon}$. The pattern $(x,x)$ could be obtained if we removed the condition, but then we would obtain unboundedly many patterns, such as $(xxx,xx)$.
Note that if we add one more sample, like $(a,a)$ to $M$ above, then we can infer $(x,x)$ as expected. ∎
As observed in Example 2.4, the result of our tuple pattern inference algorithm is not unique, but the following theorem states that our (non-deterministic) algorithm always outputs a minimal solvable pattern.
@@
The minimality property above is important for learning from positive samples only. Suppose that we are trying to learn the language described by a tuple pattern $t_{0}$. In learning from positive samples, we are given data $M$ consisting of only a (finite) subset of $\mathcal{L}(t_{0})$. Thus, our non-deterministic algorithm may output a “wrong” pattern $t_{1}$ such that $M\models t_{1}$ but $\mathcal{L}(t_{1})\neq\mathcal{L}(t_{0})$. Thanks to the theorem above, in such a case, there exists data $(s_{1},\ldots,s_{n})\in\mathcal{L}(t_{0})\setminus\mathcal{L}(t_{1})$. Thus, by adding the new positive sample $(s_{1},\ldots,s_{n})$ to $M$ and running our algorithm again, the wrong pattern $t_{1}$ will not be encountered again. Without the minimality property, it could be the case $\mathcal{L}(t_{1})\subsetneq\mathcal{L}(t_{0})$, and then the pattern $t_{1}$ cannot be refuted with only positive samples.
@@
The following theorem suggests that there exist polynomial-size data which can uniquely determine the pattern.
Theorem 2.6 (Size of Learning Data).
@@
Let $t=(p_{1},\ldots,p_{n})$ be a solvable tuple pattern such that $|p_{1}\cdots p_{n}|=m$. Then, there exists $M$ such that (i) $\mathit{size}(M)=O((m+n)\log n)$, (ii) $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}{}$ for some $\Theta$, and (iii) $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t^{\prime},\Theta^{\prime})\;\mathbin{\ \not\!\!\longrightarrow}{}$ implies $\mathcal{L}(t)=\mathcal{L}(t^{\prime})$. Furthermore, given $t$, $M$ can be constructed in polynomial time.
By Angluin’s result ([2], Theorem 1), the following is obtained as a corollary.
Corollary 2.7.
The class of solvable tuple patterns is learnable (i.e., identifiable in the limit [18]) from positive data.
The following complexity results are not directly used, but would be of interest in the field of pattern languages. Note that the membership and inclusion problems are respectively NP-complete and undecidable for both NE-pattern languages and E-pattern languages [42].
@@
    Given a solvable tuple pattern $t$ and a tuple $(s_{1},\ldots,s_{n})$ where $s_{1},\ldots,s_{n}\in\Sigma^{*}$, decide whether $(s_{1},\ldots,s_{n})\in\mathcal{L}(t)$.
@@
    Given two solvable tuple patterns $t_{1}$ and $t_{2}$, decide whether $\mathcal{L}(t_{1})\subseteq\mathcal{L}(t_{2})$.
@@
    Given two solvable tuple patterns $t_{1}$ and $t_{2}$, decide whether $\mathcal{L}(t_{1})=\mathcal{L}(t_{2})$.
@@
2.3 Extensions
We have so far considered a minimum class of solvable tuple patterns to clarify the essence. This section discusses extensions of the rewriting rules and patterns to expand the class of solvable tuple patterns. All the properties (soundness, completeness, minimality, the existence of polynomial-size data, and complexity results) remain to hold for the extensions in Sections 2.3.1 and 2.3.2; see Appendix A.2 for proofs. Further extensions of solvable tuple patterns are discussed in Section B.
2.3.1 Postfix
According to the definition of solvable tuple patterns in Section 2.2, $(x,xy)$ is solvable but $(y,xy)$ is not. To support the latter pattern, it suffices to add the following rules for $\longrightarrow$ and $\leadsto$:
$\frac{\begin{array}[]{@{}c@{}}M[*][i]\neq\widetilde{\epsilon}\quad\quad M[*][j]=\widetilde{s}\cdot M[*][i]\quad\quad x_{j}^{\prime}\mbox{ fresh}\end{array}}{\begin{array}[]{@{}c@{}}(t,[M/(x_{1},\ldots,x_{m})])\longrightarrow([x_{j}^{\prime}\cdot x_{i}/x_{j}]t,[M\{j\mapsto\widetilde{s}\}/(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{m})])\end{array}}$ (R-Postfix)
$\frac{\begin{array}[]{@{}c@{}}M[*][j]=\widetilde{s}\cdot a\quad\quad a\in\Sigma\quad\quad x_{j}^{\prime}\mbox{ fresh}\end{array}}{\begin{array}[]{@{}c@{}}(t,[M/(x_{1},\ldots,x_{m})])\longrightarrow([x_{j}^{\prime}a/x_{j}]t,[M\{j\mapsto\widetilde{s}\}/(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{m})])\end{array}}$ (R-CPostfix)
$\frac{\begin{array}[]{@{}c@{}}p_{j}=p^{\prime}_{j}\cdot p_{i}\quad\quad p_{i}\neq\epsilon\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-Postfix)
$\frac{\begin{array}[]{@{}c@{}}p_{j}=p^{\prime}_{j}\cdot a\quad\quad a\in\Sigma\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-CPostfix)
We have just replaced prefixes with postfixes.
Example 2.6.
Consider the following data:
$M=\left(\begin{array}[]{rrr}a&\ baa\\ bc&abcbc\\ \end{array}\right).$
The pattern $(x,yxx)$ is inferred as follows.
$((x_{1},x_{2}),((x_{1},x_{2}),M))$ $\longrightarrow((x_{1},x^{\prime}_{2}x_{1}),[\left(\begin{array}[]{rrr}a&\ ba\\ bc&abc\\ \end{array}\right)/(x_{1},x^{\prime}_{2})])$ (R-Postfix)
$\longrightarrow((x_{1},x^{\prime\prime}_{2}x_{1}x_{1}),[\left(\begin{array}[]{rrr}a&\ b\\ bc&a\\ \end{array}\right)/(x_{1},x^{\prime\prime}_{2})])$ $\mbox{ (\hbox{\sc R-Postfix})}.$
2.3.2 Reverse
Let us extend the pattern expressions $p$ by:
$p::=a\mid x\mid p_{1}p_{2}\mid x^{R}.$
@@ 2.3.2 Reverse
Here, $x^{R}$ represents the set of the reverse of strings represented by $x$. We extend $(\cdot)^{R}$ to the operation on patterns, by $a^{R}=a$, $(x^{R})^{R}=x$, and $(p_{1}p_{2})^{R}=p_{2}^{R}\cdot p_{1}^{R}$, so that the extended pattern language is closed under substitutions. For example, $ay/x]x^{R}=(ay)^{R}=y^{R}a$.
@@
$\frac{\begin{array}[]{@{}c@{}}M[*][i]\neq\widetilde{\epsilon}\quad\quad M[*][j]=M[*][i]^{R}\cdot\widetilde{s}\quad\quad x_{j}^{\prime}\mbox{ fresh}\end{array}}{\begin{array}[]{@{}c@{}}(t,[M/(x_{1},\ldots,x_{m})])\longrightarrow([x_{i}^{R}\cdot x_{j}^{\prime}/x_{j}]t,[M\{j\mapsto\widetilde{s}\}/(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{m})])\end{array}}$ (R-RPrefix)
$\frac{\begin{array}[]{@{}c@{}}M[*][i]\neq\widetilde{\epsilon}\quad\quad M[*][j]=\widetilde{s}\cdot M[*][i]^{R}\quad\quad x_{j}^{\prime}\mbox{ fresh}\end{array}}{\begin{array}[]{@{}c@{}}(t,[M/(x_{1},\ldots,x_{m})])\longrightarrow([x_{j}^{\prime}\cdot x_{i}^{R}/x_{j}]t,[M\{j\mapsto\widetilde{s}\}/(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{m})])\end{array}}$ (R-RPostfix)
Accordingly, we extend solvable tuple patterns by introducing the following rules:
@@
$\frac{\begin{array}[]{@{}c@{}}p_{j}=p^{\prime}_{j}\cdot p_{i}^{R}\quad\quad p_{i}\neq\epsilon\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-RPostfix)
@@
$((x_{1},x_{2},x_{3}),[M/(x_{1},x_{2},x_{3})])$
$\longrightarrow((x_{1},x_{2},x_{1}^{R}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&\ x_{2}&\ x_{3}^{\prime}\\ ab&\ cd&\ cd\\ bc&da&da\end{array}\right)$ (R-RPrefix)
$\longrightarrow((x_{1},x_{2},x_{1}^{R}x_{2}x_{3}^{\prime\prime}),\left(\begin{array}[]{rrr}x_{1}&\ x_{2}&\ x_{3}^{\prime\prime}\\ ab&\ cd&\ \epsilon\\ bc&da&\epsilon\end{array}\right)$ (R-Prefix)
$\longrightarrow((x_{1},x_{2},x_{1}^{R}x_{2}),\left(\begin{array}[]{rrr}x_{1}&\ x_{2}\\ ab&\ cd\\ bc&da\end{array}\right)$ $\mbox{ (\hbox{\sc R-Epsilon})}.$
2.3.3 Non-example: Sorting
One may also be tempted to introduce a pattern $p^{s}$, which represents the set of strings obtained by sorting strings belonging to $p$, and add the following rule:
$\frac{\begin{array}[]{@{}c@{}}M[*][i]\neq\widetilde{\epsilon}\quad\quad M[*][j]=\mathit{sort}(M[*][i])\cdot\widetilde{s}\quad\quad x_{j}^{\prime}\mbox{ fresh}\end{array}}{\begin{array}[]{@{}c@{}}(t,[M/(x_{1},\ldots,x_{m})])\longrightarrow([x_{i}^{s}\cdot x_{j}^{\prime}/x_{j}]t,[M\{j\mapsto\widetilde{s}\}/(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{m})])\end{array}}$ (R-SPrefix)
Here, $\mathit{sort}(s)$ denotes the sequence obtained by sorting $s$. We would then be able to infer a pattern $(x,x^{s})$ from data $\left(\begin{array}[]{rrr}1\cdot 3\cdot 2&\ 1\cdot 2\cdot 3\\ 5\cdot 4&4\cdot 5\end{array}\right)$.
Unfortunately, the extension above breaks the minimality property. Consider the following data:
$M=\left(\begin{array}[]{rrr}1\cdot 2&\ 2\cdot 1&\ 2\\ 3\cdot 4\cdot 5&\ 4\cdot 3\cdot 5&4\end{array}\right).$
Then, we can infer the pattern $((xy)^{s},xy,x)$ as follows.
[…]
However, the following rewriting instead infers $(z,xy,x)$, which is a superset of $((xy)^{s},xy,x)$.
[…]
The failure of minimality above can be attributed to the break of the weak confluence property $\leadsto$ (extended with the rule corresponding to R-SPrefix): $((xy)^{s},xy,x)$ is solvable because
$((xy)^{s},xy,x)\leadsto(\epsilon,xy,x)\leadsto(xy,x)\leadsto(y,x).$
However, $((xy)^{s},xy,x)\leadsto((xy)^{s},y,x)$ gets stuck.
2.3.4 Beyond Sequence Patterns

### v2

@@ 3.1 Solvability
The name “solvable tuple patterns” comes from the following fact. Suppose $t=(p_{1},\ldots,p_{n})\in\mathcal{T}_{n}$. Then, the general solution for a system of equations $p_{1}=s_{1},\ldots,p_{n}=s_{n}$ can be expressed by using (i) $s_{1},\ldots,s_{n},\epsilon$, (ii) constants $a\in\Sigma$, (iii) the (partial) operation $\ss\backslash s^{\prime}$, which is defined as $s_{0}$ such that $ss_{0}=s^{\prime}$ if $s$ is a prefix of $s^{\prime}$. and (iv) the (partial) operation $s^{\prime}/s$, which is defined as $s_{0}$ such that $s_{0}s=s^{\prime}$ if $s$ is a suffix of $s^{\prime}$. For example, $t=(x_{1}x_{2},x_{2}x_{1},x_{1})$ is solvable in the sense that the system of equations $x_{1}x_{2}=s_{1},x_{2}x_{1}=s_{2},x_{1}=s_{3}$ has a general solution $x_{1}=s_{3},x_{2}=s_{3}\backslash s_{1}$ provided $(s_{3}\backslash s_{1})\backslash s_{2}=s_{3}$ (or, equivalently $(s_{3}\backslash s_{1})\cdot s_{3}=s_{2}$). ∎
3.2 Properties of the Inference Algorithm
As stated in the following theorems, our algorithm indeed infers STPs, and is complete for the class of STPs (modulo a minor condition).
@@ 3.2 Properties of the Inference Algorithm
The following theorem, obtained as a corollary of Theorems 2.1 and 3.2, says that $\mathtt{CTPinf}(M)$ returns the least CSTP $\gamma$ such that $M\models\gamma$.
Theorem 3.3 (completeness of $\mathtt{CTPinf}$).
Let $M$ be learning data. Then $M\models\mathtt{CTPinf}(M)$, and $\mathcal{L}(\mathtt{CTPinf}(M))\subseteq\mathcal{L}(\gamma)$ for every CSTP $\gamma$ such that $M\models\gamma$.
@@
Theorem 3.2 would not hold if the assumption $M\models_{s}t$ were weakened to $M\models t$. In fact, let $t=(x,x)$ and $M=(\epsilon\ \epsilon)$. Then $M\models t$ and $t\in\mathcal{T}_{2}$, but we can only obtain $(\epsilon,\epsilon),(\epsilon,x),(x,\epsilon),(x,y)$ by rewriting. In fact, the only possible reduction sequences are:
$\Big((x_{1},x_{2}),\left(\begin{array}[]{rr}x_{1}&x_{2}\\ \epsilon&\epsilon\end{array}\right)\Big)\longrightarrow\Big((x_{1},\epsilon),\left(\begin{array}[]{r}x_{1}\\ \epsilon\end{array}\right)\Big)\longrightarrow\Big((\epsilon,\epsilon),\big(\,\big)\Big)$
$\Big((x_{1},x_{2}),\left(\begin{array}[]{rr}x_{1}&x_{2}\\ \epsilon&\epsilon\end{array}\right)\Big)\longrightarrow\Big((\epsilon,x_{2}),\left(\begin{array}[]{r}x_{2}\\ \epsilon\end{array}\right)\Big)\longrightarrow\Big((\epsilon,\epsilon),\big(\,\big)\Big).$
The rule R-Prefix is inapplicable because of the side condition $M[*][i]\neq\widetilde{\epsilon}$. The pattern $(x,x)$ could be obtained if we removed the condition, but then we would obtain unboundedly many patterns, such as $(xxx,xx)$. If we add one more sample, like $(a,a)$ to $M$ above, then we can infer $(x,x)$ as expected. ∎
As observed in Example 2.5, the output of $\mathtt{TPinf}$ is not unique due to non-determinism (whereas $\mathtt{CTPinf}$ is), but the following theorem states that $\mathtt{TPinf}$ always outputs a minimal solvable pattern.
@@
The minimality property above is important for learning from positive samples only. Suppose that we are trying to learn the language described by a tuple pattern $t_{0}$. In learning from positive samples, we are given data $M$ consisting of only a (finite) subset of $\mathcal{L}(t_{0})$. Thus, our non-deterministic algorithm may output a “wrong” pattern $t_{1}$ such that $M\models t_{1}$ but $\mathcal{L}(t_{1})\neq\mathcal{L}(t_{0})$. Thanks to the theorem above, in such a case, there exists data $\widetilde{s}\in\mathcal{L}(t_{0})\setminus\mathcal{L}(t_{1})$. Thus, by adding the new positive sample $\widetilde{s}$ to $M$ and running our algorithm again, the wrong pattern $t_{1}$ will not be encountered again. Without the minimality property, it could be the case $\mathcal{L}(t_{1})\subsetneq\mathcal{L}(t_{0})$, and then the pattern $t_{1}$ would not be refutable with only positive samples.
@@
For each STP $t$, there exist polynomial-size data that uniquely characterize $t$.
Theorem 3.5 (Characteristic Data).
@@
Let $t=(p_{1},\ldots,p_{n})$ be an STP such that $|p_{1}\cdots p_{n}|=m$. Then, there exists $M$ such that (i) $\mathit{size}(M)=O((m+n)\log n)$, (ii) for any $M^{\prime}$ such that $M\subseteq M^{\prime}\subseteq\mathcal{L}(t)$, there exists $\Theta$ such that $(\widetilde{x},[M^{\prime}/\widetilde{x}])\longrightarrow^{*}(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}{}$, and (iii) for any $M^{\prime}$ such that $M\subseteq M^{\prime}\subseteq\mathcal{L}(t)$, $(\widetilde{x},[M^{\prime}/\widetilde{x}])\longrightarrow^{*}(t^{\prime},\Theta^{\prime})\;\mathbin{\ \not\!\!\longrightarrow}{}$ implies $\mathcal{L}(t)=\mathcal{L}(t^{\prime})$. Furthermore, given $t$, $M$ can be constructed in polynomial time.
3.3 Learnability
We have so far considered an algorithm for inferring $t$ ($\gamma$, resp.) such that $M\models t$ ($M\models\gamma$, resp.), given $M$. In the whole learnig process, we need to repeatedly invoke the algorithm for gradually increasing learning data $M_{0}\subset M_{1}\subset M_{2}\subset\cdots$, until a true STP (or CSTP) is found. In Gold’s learning model [19], a class $\mathcal{C}$ of languages is learnable in the limit if there exists an algorithm $f$ such that, for any language $L\in\mathcal{C}$ and any infinite sequence $s_{0},s_{1},s_{2},\ldots$ s.t. $L=\{s_{i}\mid i\in\omega\}$, the sequence $f(\{s_{0}\}),f(\{s_{0},s_{1}\}),f(\{s_{0},s_{1},s_{2}\}),\ldots$ eventually converges to (a representation of) $L$. Both the class of STPs and that of CSTPs are learnable in this sense, and our algorithms $\mathtt{TPinf}$ and $\mathtt{CTPinf}$ serve as $f$ above, as stated below.
Theorem 3.6.
Suppose $t$ is an STP, $\mathcal{L}(t)=\{\widetilde{s}_{i}\mid i\in\omega\}$, and $M_{i}=\{\widetilde{s}_{j}\mid 0\leq j\leq i\}$. Then, there exists $k$ such that $\mathcal{L}(\mathtt{TPinf}(M_{i}))=\mathcal{L}(t)$ for all $i\geq k$. Similarly, if $\gamma$ is a CSTP, $\mathcal{L}(\gamma)=\{\widetilde{s}_{i}\mid i\in\omega\}$, and $M_{i}=\{\widetilde{s}_{j}\mid 0\leq j\leq i\}$, then there exists $k$ such that $\mathcal{L}(\mathtt{CTPinf}(M_{i}))=\mathcal{L}(\gamma)$ for all $i\geq k$.
In applications to program verification discussed later, we use CSTPs to represent and infer inductive invariants. To that end, we consider a different learning model where the goal is to find the strongest inductive invariant among those describable by a CSTP, given an oracle to check the inductiveness of a CSTP. This can be regarded as a kind of active learning framework where the oracle is a teacher. The following theorem states that our algorithm $\mathtt{CTPinf}$ can also be used in this learning framework.
Theorem 3.7.
Let $\mathcal{F}\mathbin{:}2^{\Sigma^{*}\times\cdots\times\Sigma^{*}}\to 2^{\Sigma^{*}\times\cdots\times\Sigma^{*}}$ be a monotonic function, and suppose that there exists an algorithm $g_{\mathcal{F}}$ that, given a CSTP $\gamma$, returns “None” if $\mathcal{L}(\gamma)\supseteq\mathcal{F}(\mathcal{L}(\gamma))$ holds, and returns Some(s~)Some($\widetilde${s}) for some $\widetilde{s}\in\mathcal{F}(\mathcal{L}(\gamma))\setminus\mathcal{L}(\gamma)$ if the inclusion does not hold. Then the procedure below eventually terminates and returns the least CSTP $\gamma$ such that $\mathcal{L}(\gamma)\supseteq\mathcal{F}(\mathcal{L}(\gamma))$.
$M\leftarrow\emptyset$; $\gamma\leftarrow(a,\epsilon,\ldots,\epsilon)\land(\epsilon,\epsilon,\ldots,\epsilon)$ ;
[…]
The algorithm above first sets $\gamma$ to $(a,\epsilon,\ldots,\epsilon)\land(\epsilon,\epsilon,\ldots,\epsilon)$ so that $\mathcal{L}(\gamma)=\emptyset$. It then repeatedly adds to $M$ the element returned by the oracle $g_{\mathcal{F}}$ and invokes $\mathtt{CTPinf}$. A crucial property to guarantee the theorem is that $\mathtt{CTPinf}(M)$ always returns the least CSTP $\gamma$ such that $M\subseteq\mathcal{L}(\gamma)$. See Appendix for more details.
Theorem 3.7 above would not hold if CSTPs were replaced with STPs, because there is no guarantee that there exists a least $t$ such that $\mathcal{L}(t)\supseteq\mathcal{F}(\mathcal{L}(t))$. For example, suppose that $\mathcal{F}$ is a constant function that always returns $\{(a,ab,ab),(cd,c,cde)\}$. Then, $\mathcal{L}(t_{1})\supseteq\mathcal{F}(\mathcal{L}(t_{1}))$ and $\mathcal{L}(t_{2})\supseteq\mathcal{F}(\mathcal{L}(t_{2}))$ hold for $t_{1}=(x,y,xz)$ and $t_{2}=(x,y,yz)$, but there is no STP $t$ such that $\mathcal{L}(t)\subseteq\mathcal{L}(t_{1})\cap\mathcal{L}(t_{2})$ and $\mathcal{L}(t)\supseteq\mathcal{F}(\mathcal{L}(t))$. We can, however enumerate all the minimal $t$’s such that $\mathcal{L}(t)\supseteq\mathcal{F}(\mathcal{L}(t))$.
Theorem 3.8.
[…]
3.4 Decision Problems
This section investigates decision problems on STPs. We first study the membership and inclulsion problems for STPs, which is useful for $\mathtt{CTPinf}$. Note that the membership and inclusion problems are respectively NP-complete and undecidable for both NE-pattern languages and E-pattern languages [47].
@@ 3.4 Decision Problems
    Given an STP $t$ and a tuple $\widetilde{s}\in\Sigma^{*}\times\cdots\times\Sigma^{*}$, decide whether $\widetilde{s}\in\mathcal{L}(t)$.
@@
    Given an STP $t_{2}$ and a tuple pattern $t_{1}$, decide whether $\mathcal{L}(t_{1})\subseteq\mathcal{L}(t_{2})$.
@@
    Given two STPs $t_{1}$ and $t_{2}$, decide whether $\mathcal{L}(t_{1})=\mathcal{L}(t_{2})$.
@@
We next prove the decidability of the theory of quantifier-free formulas containing formulas of the form $\widetilde{w}\in\mathcal{L}(t)$. This decidability is important for constructing the oracle $g_{\mathcal{F}}$ in Theorem 3.7, and plays an important role in the applications to program verification discussed in Section 5.
Definition 3.1.
The set of (quantifier-free) STP formulas, ranged over by $\varphi$, is given by:
$\varphi::=w_{1}=w_{2}\mid w_{1}\neq w_{2}\mid\widetilde{w}\in\mathcal{L}(t)\mid\widetilde{w}\not\in\mathcal{L}(t)\mid\varphi_{1}\land\varphi_{2}\mid\varphi_{1}\lor\varphi_{2}\qquad w\in(\Sigma\cup\mathcal{V})^{*}.$
Here, $t$ ranges over the set of STPs.
For example, $(x,y,z)\in\mathcal{L}(t)\land w=yz\land(xy,y,w)\not\in\mathcal{L}(t)$ (which is the negation of $(x,y,z)\in\mathcal{L}(t)\land w=yz$ implies $(xy,y,w)\in\mathcal{L}(t)$”), where $t=(u,v,uv)$ is an STP formula. The semantics of STP formulas is defined in an obvious manner, and is therefore deferred to the Appendix. The only point to note is that the variables occurring in $t$ are implicitly bound within $\mathcal{L}(t)$; for exapmle, $(x,y,z)\in\mathcal{L}(u,v,uv)$ is equivalent to $(x,y,z)\in\mathcal{L}(y,x,yx)$ (although we usually avoid variable overloading to prevent confusion).
Remark 3.5.
The definition of quantifier-free STP formulas above can be actually simplified to:
$\varphi::=\widetilde{w}\in\mathcal{L}(t)\mid\widetilde{w}\not\in\mathcal{L}(t)\mid\varphi_{1}\land\varphi_{2}\mid\varphi_{1}\lor\varphi_{2},$
because $w_{1}=w_{2}$ and $w_{1}\neq w_{2}$ can be expressed by $(w_{1},w_{2})\in\mathcal{L}(x,x)$ and $(w_{1},w_{2})\not\in\mathcal{L}(x,x)$ respectively. ∎
The satisfiability of STP formulas is decidable, by reduction to that of word equations [41, 51].
Theorem 3.10.
Given an STP formula $\varphi$, one can effectively construct an equi-satisfiable STP formula $\varphi^{\prime}$ that contains no subformulas of the form $w_{1}\neq w_{2}$, $(w_{1},\ldots,w_{k})\in\mathcal{L}(t)$, or $(w_{1},\ldots,w_{k})\not\in\mathcal{L}(t)$. Therefore, the satisfiability of STP formulas is decidable.
Encoding of a formula $\widetilde{w}\not\in\mathcal{L}(t)$ into word equations relies on the solvability of $t$. In fact, if we allowed arbitrary tuple patterns, then the satisfiability would become undecidable, because $\exists x\in\Sigma^{*}.(x\in\mathcal{L}(p_{1})\land x\not\in\mathcal{L}(p_{2}))$ if and only if $\mathcal{L}(p_{1}){\not\subseteq}\mathcal{L}(p_{2})$ and the latter is undecidable [28, 47].
4 Extensions and Variations of Solvable Tuple Patterns
This section discusses extensions and variations of STPs. Section 4.1 extends STPs with the reverse pattern $x^{R}$, which is needed to handle the $\mathit{Reva}$ example in Section 1. Section 4.2 discusses variations of STPs to reason about sets and multisets.
4.1 Reverse
@@
Let us extend the set of pattern expressions (ranged over by $p$) to $(\Sigma\cup\mathcal{V}\cup\mathcal{V}^{R})^{*}$, where $\mathcal{V}^{R}=\{x^{R}\mid x\in\mathcal{V}\}$. Here, $x^{R}$ represents the set of the reverse of strings represented by $x$. We extend $(\cdot)^{R}$ to the operation on patterns, by $a^{R}=a$, $(x^{R})^{R}=x$, and $(p_{1}p_{2})^{R}=p_{2}^{R}\cdot p_{1}^{R}$, so that the set of extended pattern expressions is closed under substitutions. For example, $ay/x]x^{R}=(ay)^{R}=y^{R}a$.
@@
$\frac{\begin{array}[]{@{}c@{}}M[*][i]\neq\widetilde{\epsilon}\quad\quad M[*][j]=M[*][i]^{R}\cdot\widetilde{s}\quad\quad x_{j}^{\prime}\mbox{ fresh}\quad\quad\widetilde{x}=(x_{1},\ldots,x_{m})\end{array}}{\begin{array}[]{@{}c@{}}(t,[M/\widetilde{x}])\longrightarrow([x_{i}^{R}\cdot x_{j}^{\prime}/x_{j}]t,[M\{j\mapsto\widetilde{s}\}/(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{m})])\end{array}}$ (R-RPrefix)
$\frac{\begin{array}[]{@{}c@{}}M[*][i]\neq\widetilde{\epsilon}\quad\quad M[*][j]=\widetilde{s}\cdot M[*][i]^{R}\quad\quad x_{j}^{\prime}\mbox{ fresh}\quad\quad\widetilde{x}=(x_{1},\ldots,x_{m})\end{array}}{\begin{array}[]{@{}c@{}}(t,[M/\widetilde{x}])\longrightarrow([x_{j}^{\prime}\cdot x_{i}^{R}/x_{j}]t,[M\{j\mapsto\widetilde{s}\}/(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{m})])\end{array}}$ (R-RSuffix)
Here, $M[*][i]^{R}$ denotes the pointwise extension of the reverse operation. Accordingly, we extend solvable tuple patterns by adding the following rules:
@@
$\frac{\begin{array}[]{@{}c@{}}p_{j}=p^{\prime}_{j}\cdot p_{i}^{R}\quad\quad p_{i}\neq\epsilon\end{array}}{\begin{array}[]{@{}c@{}}(p_{1},\ldots,p_{n})\leadsto(p_{1},\ldots,p_{j-1},p^{\prime}_{j},p_{j+1},\ldots,p_{n})\end{array}}$ (PR-RSuffix)
@@
$\Big((x_{1},x_{2},x_{3}),\left(\begin{array}[]{rrr}x_{1}&x_{2}&x_{3}\\ ab&\ cd&\ bacd\\ bc&da&cbda\end{array}\right)\Big)$
$\longrightarrow\Big((x_{1},x_{2},x_{1}^{R}x_{3}^{\prime}),\left(\begin{array}[]{rrr}x_{1}&\ x_{2}&\ x_{3}^{\prime}\\ ab&\ cd&\ cd\\ bc&da&da\end{array}\right)\Big)$ (R-RPrefix)
$\longrightarrow\Big((x_{1},x_{2},x_{1}^{R}x_{2}x_{3}^{\prime\prime}),\left(\begin{array}[]{rrr}x_{1}&\ x_{2}&\ x_{3}^{\prime\prime}\\ ab&\ cd&\ \epsilon\\ bc&da&\epsilon\end{array}\right)\Big)$ (R-Prefix)
$\longrightarrow\Big((x_{1},x_{2},x_{1}^{R}x_{2}),\left(\begin{array}[]{rrr}x_{1}&\ x_{2}\\ ab&\ cd\\ bc&da\end{array}\right)\Big)$ $\mbox{ (\hbox{\sc R-Epsilon})}.$
Remark 4.1.
One may also be tempted to extend patterns with $x^{S}$, which represents the sequence obtained by sorting $x$. This is problematic, however, because $(\cdot)^{S}$ does not commute with concatenation: $(y\cdot z)^{S}\neq y^{S}\cdot z^{S}$. One way to reason about sorted sequences is to introduce a new class of patterns called sort patterns, given by: $p_{S}::=x\mid u$ and $u::=x^{S}\mid u_{1}\star u_{2}$. Here, $u_{1}\star u_{2}$ denotes the set of sorted sequences obtained by merging sorted sequences represented by $u_{1}$ and $u_{2}$. Sort patterns can naturally be extended to tuple sort patterns. For example, $(x,x^{S})$ denotes the set of pairs, whose second element is obtained by sorting the first element, and $(x,y^{S},x^{S}\star y^{S})$ denotes the set of triples $(s_{1},s_{2},s_{3})$, where $s_{3}$ is obtained by sorting $s_{1}$ and merging it with $s_{2}$. We can then design an inference algorithm for (solvable) tuple sort patterns in a manner similar to ordinary tuple patterns. The details will be described in a separate paper. ∎
4.2 Beyond Sequence Patterns